# Deep Research Agents: A Systematic Examination And Roadmap

### v1

@@ 3 Deep Research: Search Engine, Tool Use, Workflow, Tuning, Non-parametric Continual Learning
Browser-based retrieval provides DR agents with dynamic, flexible, and interactive access to multimodal and unstructured web content through simulated human-like browser interactions. For example, Manus AI’s browsing agent operates a sandboxed Chromium instance for each research session, programmatically opening new tabs, issuing search queries, clicking through result links, scrolling pages until content thresholds are met, filling out form elements when necessary, executing in-page JavaScript to reveal lazily loaded sections, and downloading files or PDFs for local analysis manus2025 . Although OpenAI DR, Grok DeepSearch, and Gemini 2.5 DR do not publicly disclose the implementation details of their browsing capabilities, their ability to handle interactive widgets, dynamically rendered content, and multi-step navigation strongly suggests that they too employ comparable headless-browser frameworks behind the scenes. Among open-source studies, AutoAgent yu2024auto operates within a BrowserGym environment to scroll, interact with page components, and download files when APIs are unavailable yu2024auto ; DeepResearcher zheng2025deepresearcherscalingdeepresearch employs a dedicated Web Browsing Agent that, upon receiving a browse request, processes each segment of a webpage in turn, decides whether to continue to subsequent segments based on relevance, and incrementally aggregates pertinent information into a short-term memory buffer before returning it for reasoning. While browser-based retrieval excels at capturing real-time and deeply nested content that API calls cannot reach, it also incurs greater latency, resource consumption, and complexity in handling page variability and errors, suggesting that DR agents may benefit from hybrid architectures that combine the efficiency of API-based methods with the comprehensiveness of browser-driven exploration.
@@ 3.2 Tool Use: Empowering Agents with Extended Functionalities
By integrating data analytics modules, DR agents transform raw retrievals into structured insights by computing summary statistics, generating interactive visualizations and conducting quantitative model evaluations, thereby accelerating hypothesis testing and decision-making. Many commercial DR agents have implemented analytics features such as charting, table generation and statistical analysis, either locally or via remote services. However, most of these systems have not publicly disclosed technical details of their implementations. In contrast, academic studies often provide concrete examples: CoSearchAgent integrates SQL-based queries within team communication platforms to run aggregate analyses and produce reports; AutoGLM extracts and analyzes structured datasets directly from table-based web interfaces; and Search-o1’s Reason-in-Documents component refines lengthy retrieved texts before extracting key metrics for downstream evaluation.
@@
Multimodal processing and generation tools enable DR agents to integrate, analyze and generate heterogeneous data such as text, images, audio and video within a unified reasoning pipeline, thereby enriching their contextual understanding and broadening the range of their outputs. Only a subset of mature commercial and open-source projects, for example Manus, OWL, AutoAgent, AutoGLM, OpenAI, Gemini, Perplexity and Grok DeepSearch, support this capability, whereas most academic prototypes have not implemented it, often due to the high computational cost. As the typical open source studies, OWL and Openmanus extend their pipelines to include interactions with platforms such as GitHub, Notion and Google Maps and to leverage numerical libraries such as Sympy and Excel for combined data analysis and multimodal media processing owl2025 ; openmanus2025 .
@@
Most recently, the boundaries of DR agents have been progressively expanded through integrating computer-assisted task execution capabilities (i.e., computer use). For example, Zhipu AI introduced AutoGLM Rumination zhipu2025autoglm , a RL-based system incorporating self-reflection and iterative refinement mechanisms, which significantly enhances multi-step reasoning and advanced function-calling abilities. Specifically, AutoGLM Rumination autonomously interacts with web environments, executes code, invokes external APIs, and effectively accomplishes sophisticated tasks, including data retrieval, analysis, and structured generation of comprehensive reports. Comparison with OpenAI’s DR: While OpenAI DR primarily focus on intricate reasoning and information retrieval, AutoGLM Rumination exhibits superior autonomy in practical execution. This enhanced autonomy allows it to transform abstract analytical insights into concrete operational tasks, such as automated interactions with web interfaces and real-time data processing. Moreover, AutoGLM Rumination addresses and resolves limitations inherent in simulated browsing environments by seamlessly integrating advanced reasoning capabilities with authentic browser-based interactions. Therefore, the agent gains reliable access to user-authenticated resources, including platforms such as CNKI, Xiaohongshu, and WeChat official accounts. Such integration significantly elevates the agent’s autonomy and adaptability in both information acquisition and execution of real-world tasks.
@@ 3.3 Architecture and Workflow
As shown in Figure 4, this section systematically analyzes the construction of DR systems, focusing on workflows categorized into static and dynamic types. We then discuss planning strategies, which enhance task allocation and execution through three distinctive user interaction types to clarify intent: planning-only (direct planning without clarifying user intent), intent-to-planning (clarifying intent before planning to align the task with user goals), and unified intent-planning (generating a plan and requesting user confirmation). The distinction between single-agent and multi-agent systems is examined in the context of dynamic workflows, emphasizing specialization in task management. Additionally, we examine memory mechanisms for managing and integrating retrieved information, which enhance the performance and adaptability of DR systems.
@@
Figure 4: Comparison of DR Workflows: (1) Static vs. Dynamic Workflows: Static workflows rely on predefined task sequences, while dynamic workflows allow LLM-based task planning. (2) Planning Strategies: Three types include: planning-only (direct planning without clarifying user intent), intent-to-planning (clarifying intent before planning), and unified intent-planning (generating a plan and requesting user confirmation). (3) Single-Agent vs. Multi-Agent: Dynamic workflows can be categorized to dynamic-multi-agent systems (tasks distributed across specialized agents) or dynamic-single-agent systems (a LRM autonomously updates and executes tasks).
@@
Static workflows rely on manually predefined task pipelines, decomposing research processes into sequential subtasks executed by dedicated agents. These workflows follow explicitly structured procedures, making them particularly suitable for well-defined, structured research scenarios. For instance, AI Scientist lu2024aiscientistfullyautomated automates scientific discovery through distinct sequential phases, including ideation, experimentation, and reporting. Similarly, Agent Laboratory schmidgall2025agent segments research activities into formalized stages, such as literature review, experimentation, and synthesis of findings. Extending this static paradigm further, AgentRxiv schmidgall2025agentrxiv incorporates inter-agent collaboration mechanisms, enabling incremental knowledge reuse through sharing intermediate research outcomes among specialized agents. Whist their ease of implementation and structured clarity, static workflows suffer from limited generalization capabilities, as each distinct task necessitates a specifically tailored pipeline.
@@ 3.3.2 Dynamic Workflows: Planning Strategies
    The Planning-Only approach directly generates task plans based solely on initial user prompts without actively engaging in further clarification, adopted by the majority of existing DR agents, including Grok grokdeepresearch , H2O h2oai and Manus manus2025 .
@@
    The Intent-to-Planning strategy actively clarifies user intent prior to planning through targeted questions, subsequently generating tailored task sequences based on clarified user inputs; this method is utilized by OpenAI DR openai2025deepresearch .
@@
    The Unified Intent-Planning approach synthesizes these methods by generating a preliminary plan from the initial prompt, together with interactively engaging the user to confirm or revise the proposed plan. Gemini DR geminideepresearch is representative of this strategy, effectively adopts the strength of user-guided refinement.
@@
Dynamic workflows of DR agents can be differentiated based on agent architectures into single-agent and multi-agent frameworks, each exhibiting distinct characteristics concerning task specialization, coordination complexity, and scalability of execution.
@@
Dynamic single-agent systems integrate planning, tool invocation, and execution within a unified LRM, streamlining task management into a cohesive cognitive loop. Single-agent architectures autonomously refine task plans and invoke appropriate tools based on evolving contexts, typically without explicit inter-agent coordination. Compared to multi-agent architectures, single-agent systems enable direct end-to-end reinforcement learning (RL) optimization across the entire workflow, facilitating smoother and more coherent integration of reasoning, planning, and tool invocation. Systems such as Agent-R1 Agent-R1 , ReSearch chen2025learning , and Search-R1 jin2025search exemplify this paradigm through iterative cycles of explicit reasoning, action, and reflection, aligning with the ReAct framework yao2023reactsynergizingreasoningacting . However, this streamlined approach places significant demands on the foundation model’s reasoning capabilities, contextual understanding, and autonomous selection and invocation of tools. Additionally, the tightly integrated nature of single-agent systems may limit modular flexibility, complicating independent scaling or optimization of individual functional components.
@@
Dynamic multi-agent systems leverage multiple specialized agents to collaboratively execute subtasks generated and dynamically allocated through adaptive planning strategies. These systems typically employ hierarchical or centralized planning mechanisms, wherein a coordinator agent continuously assigns and redistributes tasks based on real-time feedback and replanning. Representative frameworks include OpenManus openmanus2025 and Manus manus2025 , both adopting hierarchical planner-toolcaller architectures. Similarly, OWL owl2025 includes a workforce-oriented model, utilizing a central manager agent to orchestrate task distribution among specialized execution agents. Furthermore, Alita qiu2025alita incorporates a self-evolution mechanism into DR agents, allowing the agent to online instantiate and configure new MCP servers tailored to specific tasks and environmental conditions. Such multi-agent configurations effectively handle complex, parallelizable research tasks, thereby enhancing flexibility and scalability in open-ended research scenarios. Nevertheless, a major current challenge of multi-agent systems lies in the inherent complexity of coordinating multiple independent agents, making it difficult to conduct effective end-to-end reinforcement learning optimization.
3.3.4 Memory Mechanism for Long-Context Optimization

### v2

@@ 3 Deep Research: Search Engine, Tool Use, Workflow, Tuning, Non-parametric Continual Learning
Browser-based retrieval provides DR agents with dynamic, flexible, and interactive access to multimodal and unstructured web content through simulated human-like browser interactions. For example, Manus AI’s browsing agent operates a sandboxed Chromium instance for each research session, programmatically opening new tabs, issuing search queries, clicking through result links, scrolling pages until content thresholds are met, filling out form elements when necessary, executing in-page JavaScript to reveal lazily loaded sections, and downloading files or PDFs for local analysis (manus2025, ). Although OpenAI DR, Grok DeepSearch, and Gemini 2.5 DR do not publicly disclose the implementation details of their browsing capabilities, their ability to handle interactive widgets, dynamically rendered content, and multi-step navigation strongly suggests that they too employ comparable headless-browser frameworks behind the scenes. Among open-source studies, AutoAgent (yu2024auto, ) operates within a BrowserGym environment to scroll, interact with page components, and download files when APIs are unavailable; DeepResearcher (zheng2025deepresearcherscalingdeepresearch, ) employs a dedicated Web Browsing Agent that, upon receiving a browse request, processes each segment of a webpage in turn, decides whether to continue to subsequent segments based on relevance, and incrementally aggregates pertinent information into a short-term memory buffer before returning it for reasoning. Kimi-Researcher (kimi2025researcher, ) uses an internal search engine with a text-based browser to retrieve information. Search-R1 and MiroRL (jin2025search, ; 2025mirorl, ) employ both search and browser tools during training. AutoGLM (zhipu2025autoglm, ) operationalises browsing through a plan–execute loop that opens and reads web pages, layering “rumination” cycles over browser actions to refine evidence and produce long-form reports. Genspark Super Agent (genspark, ) orchestrates a mixture-of-agents in which a research sub-agent performs web search and page reading, handing structured notes to downstream writing/analysis agents rather than relying on a single monolithic browser. SimpleDeepSearcher (SimpleDeepSearcher, ) follows a lightweight search-fetch-summarise loop using web search APIs plus HTTP fetching in lieu of full browser automation, caching pages and compressing them before further reasoning. Tool-Star (dong2025tool, ) explicitly separates a Search Engine tool from a Web Browser Agent: after link retrieval, the browser agent opens pages, extracts salient snippets, and returns compressed evidence to the planner. AgenticSeek (agenticseek, ) couples a local metasearch front end with a headless, stealth browser so the agent can click, scroll, and submit forms on live sites, exposing knobs for budget and anti-bot robustness. AWorld (aworld2025, ) provides a multi-agent runtime with built-in browser automation and tracing, enabling teams of agents to divide research, browsing, and synthesis for deep-research workflows over dynamic sites. WebThinker (Li2025webthinker, ) performs information seeking by issuing searches and following links on returned result pages. WebDancer, WebSailor, and WebShaper (wu2025webdancer, ; li2025websailor, ; tao2025webshaper, ) combine web search with on-page navigation as a minimal sufficient toolset, linking structured tool invocations with observations in a ReAct-style closed loop: they first locate candidate sources horizontally, then drill down vertically, trading a compact action space for training stability and stronger generalisation. WebWatcher (geng2025webwatcher, ) uses the Google SerpApi for multimodal search and applies OCR-based image processing to emulate browser interactions. While browser-based retrieval excels at capturing real-time and deeply nested content that API calls cannot reach, it also incurs greater latency, resource consumption, and complexity in handling page variability and errors, suggesting that DR agents may benefit from hybrid architectures that combine the efficiency of API-based methods with the comprehensiveness of browser-driven exploration.
@@ 3.2 Tool Use: Empowering Agents with Extended Functionalities
O-agents (zhu2025oagents, ) Jun-2025
Kimi-Researcher (kimi2025researcher, ) Jun-2025
Agent-KB (tang2025agent, ) Jul-2025
AWorld (aworld2025, ) Jul-2025
Cognitive Kernel-Pro (wan2025cognitive, ) Aug-2025
WebWatcher (geng2025webwatcher, ) Aug-2025
MiroRL (2025mirorl, ) Aug-2025
@@
By integrating data analytics modules, DR agents transform raw retrievals into structured insights by computing summary statistics, generating interactive visualisations and conducting quantitative model evaluations, thereby accelerating hypothesis testing and decision-making. Many commercial DR agents have implemented analytics features such as charting, table generation and statistical analysis, either locally or via remote services. However, most of these systems have not publicly disclosed technical details of their implementations. In contrast, academic studies often provide concrete examples: CoSearchAgent (gong2024cosearchagent, ) integrates SQL-based queries within team communication platforms to run aggregate analyses and produce reports; AutoGLM (zhipu2025autoglm, ) extracts and analyses structured datasets directly from table-based web interfaces.
@@
Multimodal processing and generation tools enable DR agents to integrate, analyse and generate heterogeneous data such as text, images, audio and video within a unified reasoning pipeline, thereby enriching their contextual understanding and broadening the range of their outputs. Only a subset of mature commercial and open-source projects, for example Manus (manus2025, ), OWL (owl2025, ), AutoAgent (tang2025autoagentfullyautomatedzerocodeframework, ), AutoGLM (zhipu2025autoglm, ), OpenAI (openai2025deepresearch, ), Gemini (geminideepresearch, ), Perplexity (perplexitydeepresearch, ) and Grok DeepSearch (grokdeepresearch, ), support this capability, whereas most academic prototypes have not implemented it, often due to the high computational cost. As the typical open source studies, OWL and Openmanus extend their pipelines to include interactions with platforms such as GitHub, Notion and Google Maps and to leverage numerical libraries such as Sympy and Excel for combined data analysis and multimodal media processing (owl2025, ; openmanus2025, ).
@@
Most recently, the boundaries of DR agents have been progressively expanded through integrating computer-assisted task execution capabilities (i.e., computer use). For example, Zhipu AI introduced AutoGLM Rumination (zhipu2025autoglm, ), an RL-based system incorporating self-reflection and iterative refinement mechanisms, which significantly enhances multi-step reasoning and advanced function-calling abilities. Specifically, AutoGLM Rumination (zhipu2025autoglm, ) autonomously interacts with web environments, executes code, invokes external APIs, and effectively accomplishes sophisticated tasks, including data retrieval, analysis, and structured generation of comprehensive reports. Comparison with OpenAI’s DR: While OpenAI DR primarily focus on intricate reasoning and information retrieval, AutoGLM Rumination exhibits superior autonomy in practical execution. This enhanced autonomy allows it to transform abstract analytical insights into concrete operational tasks, such as automated interactions with web interfaces and real-time data processing. Moreover, AutoGLM Rumination addresses and resolves limitations inherent in simulated browsing environments by seamlessly integrating advanced reasoning capabilities with authentic browser-based interactions. Therefore, the agent gains reliable access to user-authenticated resources, including platforms such as CNKI, Xiaohongshu, and WeChat official accounts. Such integration significantly elevates the agent’s autonomy and adaptability in both information acquisition and execution of real-world tasks.
OpenAI DR, Perplexity DR, Grok DR, H2O, Manus, Genspark Super Agent openai2025deepresearch ; perplexitydeepresearch ; grokdeepresearch ; h2oai ; manus2025 ; genspark orchestrate multi-step web research with browser or tool use, code execution, and analytics to produce cited, structured reports at industrial scale. OWL, OpenManus, Suna, DeerFlow, WebThinker, AgenticSeek, and AWorld (owl2025, ; openmanus2025, ; sunaai, ; deerflow, ; Li2025webthinker, ; agenticseek, ; aworld2025, ) provide open-source stacks that integrate browser automation, code interpreters, and (often) MCP-style tooling for end-to-end deep research workflows. AutoGLM Rumination, Tool-Star, Kimi-Researcher, and MiroRL (zhipu2025autoglm, ; dong2025tool, ; kimi2025researcher, ; 2025mirorl, ) use reinforcement learning or self-reflection to plan searches, invoke tools and code, and improve multi-step reasoning autonomy. The AI Scientist, Storm, Agent Laboratory, Agent-R1, AutoAgent, and CoSearchAgent (lu2024aiscientistfullyautomated, ; shao2024assistingwritingwikipedialikearticles, ; schmidgall2025agent, ; Agent-R1, ; tang2025autoagentfullyautomatedzerocodeframework, ; gong2024cosearchagent, ) automate literature review, code/experiment execution, and structured drafting through tool-augmented research pipelines. Towards an AI co-scientist and O-agents (gottweis2025towards, ; zhu2025oagents, ) outline blueprints and empirical recipes for building, training, and evaluating effective tool-using research agents. Agent-KB (tang2025agent, ) introduces a KB-driven framework for cross-domain experience transfer to boost generalisation on complex tasks. Microsoft Copilot Researcher (microsoft_copilot_researcher, ) embeds multi-step research and analytics in the Microsoft 365 ecosystem, producing reports and charts within enterprise workflows. Alita (qiu2025alita, ) explores self-evolving agents that generate and wrap MCP tools, alongside code execution, to extend capabilities with minimal predefined schemas.
@@ 3.3 Architecture and Workflow
As shown in Figure 4, this section systematically analyses the construction of DR systems, focusing on workflows categorised into static and dynamic types. We first introduce the static workflows and then discuss planning strategies, which enhance task allocation and execution through three distinctive user interaction types to clarify intent: planning-only (direct planning without clarifying user intent), intent-to-planning (clarifying intent before planning to align the task with user goals), and unified intent-planning (generating a plan and requesting user confirmation). The distinction between single-agent and multi-agent systems is examined in the context of dynamic workflows, emphasising specialisation in task management. Additionally, we examine memory mechanisms for managing and integrating retrieved information, which enhance the performance and adaptability of DR systems.
@@
Figure 4: Comparison of DR Workflows: (1) Static vs. Dynamic Workflows: Static workflows rely on predefined task sequences, while dynamic workflows allow LLM-based task planning. (2) Planning Strategies: Three types include: planning-only (direct planning without clarifying user intent), intent-to-planning (clarifying intent before planning), and unified intent-planning (generating a plan and requesting user confirmation). (3) Single-Agent vs. Multi-Agent: Dynamic workflows can be categorised into dynamic-multi-agent systems (tasks distributed across specialised agents) or dynamic-single-agent systems (an LRM autonomously updates and executes tasks). For a continuously updated, per-work taxonomy of DR workflows, see https://github.com/ai-agents-2030/awesome-deep-research-agent.
@@
Static workflows rely on manually predefined task pipelines, decomposing research processes into sequential subtasks executed by dedicated agents. These workflows follow explicitly structured procedures, making them particularly suitable for well-defined, structured research scenarios. For instance, AI Scientist (lu2024aiscientistfullyautomated, ) automates scientific discovery through distinct sequential phases, including ideation, experimentation, and reporting. Similarly, Agent Laboratory (schmidgall2025agent, ) segments research activities into formalised stages, such as literature review, experimentation, and synthesis of findings. Extending this static paradigm further, AgentRxiv (schmidgall2025agentrxiv, ) incorporates inter-agent collaboration mechanisms, enabling incremental knowledge reuse through sharing intermediate research outcomes among specialised agents.
Whist their ease of implementation and structured clarity, static workflows suffer from limited generalisation capabilities, as each distinct task necessitates a specifically tailored pipeline.
@@ 3.3.2 Dynamic Workflows: Planning Strategies
    The Planning-Only approach directly generates task plans based solely on initial user prompts without actively engaging in further clarification, adopted by the majority of existing DR agents, including Grok (grokdeepresearch, ), H2O (h2oai, ) and Manus (manus2025, ).
@@
    The Intent-to-Planning strategy actively clarifies user intent prior to planning through targeted questions, subsequently generating tailored task sequences based on clarified user inputs; this method is utilised by OpenAI DR (openai2025deepresearch, ).
@@
    The Unified Intent-Planning approach synthesises these methods by generating a preliminary plan from the initial prompt, together with interactively engaging the user to confirm or revise the proposed plan. Gemini DR (geminideepresearch, ) is representative of this strategy, effectively adopts the strength of user-guided refinement.
@@
Dynamic workflows of DR agents can be differentiated based on agent architectures into single-agent and multi-agent frameworks, each exhibiting distinct characteristics concerning task specialisation, coordination complexity, and scalability of execution.
@@
Dynamic single-agent systems integrate planning, tool invocation, and execution within a unified LRM, streamlining task management into a cohesive cognitive loop. Single-agent architectures autonomously refine task plans and invoke appropriate tools based on evolving contexts, typically without explicit inter-agent coordination. Compared to multi-agent architectures, single-agent systems enable direct end-to-end reinforcement learning (RL) optimisation across the entire workflow, facilitating smoother and more coherent integration of reasoning, planning, and tool invocation. Systems such as Search-o1 (li2025search, ), R1-Searcher (song2025r1, ), DeepResearcher (zheng2025deepresearcherscalingdeepresearch, ), WebDancer (wu2025webdancer, ), WebSailor (li2025websailor, ), PANGU Deepdiver (shi2025pangu, ), Agent-R1 (Agent-R1, ), ReSearch (chen2025learning, ), Search-R1 (jin2025search, ), WebWatcher (wu2025webwalker, ), MiroRL (2025mirorl, ), Memento (zhou2025agentfly, ) and Kimi-Researcher (kimi2025researcher, ) exemplify this paradigm through iterative cycles of explicit reasoning, action, and reflection, aligning with the ReAct framework (yao2023reactsynergizingreasoningacting, ). However, this streamlined approach places significant demands on the foundation model’s reasoning capabilities, contextual understanding, and autonomous selection and invocation of tools. Additionally, the tightly integrated nature of single-agent systems may limit modular flexibility, complicating independent scaling or optimisation of individual functional components.
@@
Dynamic multi-agent systems leverage multiple specialised agents to collaboratively execute subtasks generated and dynamically allocated through adaptive planning strategies. These systems typically employ hierarchical or centralised planning mechanisms, wherein a coordinator agent continuously assigns and redistributes tasks based on real-time feedback and replanning. Representative frameworks include OpenManus (openmanus2025, ) and Manus (manus2025, ), both adopting hierarchical planner-toolcaller architectures. Similarly, OWL (owl2025, ) includes a workforce-oriented model, utilising a central manager agent to orchestrate task distribution among specialised execution agents. Furthermore, Alita (qiu2025alita, ) incorporates a self-evolution mechanism into DR agents, allowing the agent to online instantiate and configure new MCP servers tailored to specific tasks and environmental conditions. AWorld (aworld2025, ) is an open-source framework for building, orchestrating, and training tools using agents and larger multi-agent systems, offering memory and context services and MCP tool integration for scalable evaluation and self-improvement. Webwalker (wu2025webwalker, ) mimics human-like web navigation through an explore-critic paradigm. WebThinker (Li2025webthinker, ) uses both executed and auxiliary agents to autonomously search, deeply explore web pages, and draft research reports. Such multi-agent configurations effectively handle complex, parallelizable research tasks, thereby enhancing flexibility and scalability in open-ended research scenarios. Nevertheless, a major current challenge of multi-agent systems lies in the inherent complexity of coordinating multiple independent agents, making it difficult to conduct effective end-to-end reinforcement learning optimisation.
3.3.4 Memory Mechanism for Long-Context Optimisation